The area of a rectangle is 93 square meters. If the length is 2 times the width, then find the dimensions of the rectangle. Round off your answers to the nearest hundredth.
Answer by Herdaysoji(25)

area is 93=length*width
length=2*width
93=2*width*width
93=2(width)^2
so, 93/2=width^2
46.5=width^2
width=root 46.5
width=6.8190.....
round off to 6.82
length=2*6.82=13.64